Find the domain of g(t) = 5^t
Answer by gonzo(654)

the domain of the function is t.
-----
t can be any value, so the domain of the function is the set of all real numbers.
g(t) can also be any value, so the range of the function is the set of all real numbers.
-----
t can be 0:
t(0) = 5^0 = 1
-----
t can be any positive number:
t(5.4) = 5^(5.4) = 5948.918...
-----
t can be any negative number:
5(-5.4) = 5^(-5.4) = .000168.....
-----